[|111||230||383|] ,is the matrix invertible?
1
Expert's answer
2020-05-06T20:07:27-0400

A matrix is invertible if its determinant is non-zero.

We will check the determinant of the given matrix:

"\begin{vmatrix}
 1 & 1 & 1 \\
 2 & 3 & 0 \\
 3 & 8 & 3 \\
\end{vmatrix}" = 1"\begin{vmatrix}
 3 & 0 \\
 8 & 3
\end{vmatrix}" -1"\begin{vmatrix}
 2 & 0 \\
 3 & 3
\end{vmatrix}" +1"\begin{vmatrix}
 2 & 3 \\
 3 & 8
\end{vmatrix}" = 1(9)-1(6)+1(7) = 10 which is not equal to zero.

Hence, the given matrix is invertible.